# ViMultiChoice: Toward a Method That Gives Explanation for Multiple-Choice Reading Comprehension in Vietnamese

### v1

@@ 5.3 Configuration
All baseline models and the proposed method ViMultiChoice were trained using the Adam optimizer Kingma and Ba (2014). For baseline models based on static word embeddings, we employed pretrained word embedding PhoW2V Nguyen et al. (2020a) at syllable level, while the remaining baselines utilized the pretrained language model ViT5 Phan et al. (2022) to encode input sequences. All methods were trained with a learning rate of $5\times 10^{-5}$, a batch size of 32 on an NVIDIA A100 GPU. The interruption of the training process is evolve if there is no improvement in F1-macro after 10 consecutive epochs.
Moreover, we deployed an information retriever to obtain the discrete sentences for training baselines and ViMultiChoice method on the ViRCSoSciD dataset. Through our experiments, we found that VietBGE, the pretrained BGE-M3 Chen et al. (2024) specially fine-tuned for Vietnamese is the most effective retriever.
@@ 6 Results
6.1 Results of LLMs on the ViRCSoSciD Dataset
# Dataset ViWordFormer Training Type Multiple Choice Explanation
Accuracy F1-macro BLEU-4 ROUGE-L
1 ViRCSoSciD Single Task 61.81 61.79 - -
2 62.98 63.03 - -
3 ViMMRC 2.0 56.27 56.29 - -
4 60.32 60.23 - -
5 ViRCSoSciD Multitask 62.51 62.51 20.23 44.71
6 64.38 64.38 34.85 56.78
Table 4: Ablation study of ViWordFormer module for ViMultiChoice on the ViRCSoSciD and ViMMRC 2.0 datasets.
Table 1 denotes the results of LLMs for the multiple choice task while Table 2 indicates the results of LLMs for explanation task on our ViRCSoSciD dataset. Mitral-Large obtained the best performance on F1-macro on multiple-choice task and BLEU-4 as well as ROUGE-L on explanation generator. Gemeni-Flash achieved the second position on both tasks, especially it has the best accuracy score on giving options for multiple-choice.
However, the performance of other LLMs is left far away from the two best LLMs although they were pretrained on extremely huge corpora. Moreover, Mistral-Large still gives a sloghtly low ROUGE-L score for explanation task. These numbers highlights the challenging of our dataset as well as the need of developing more effective LLM particularly for Vietnamese and other low-resource languages.
6.2 Results of ViMultiChoice
@@
Especially, in comparison to the original OCN, our ViMultiChoice can easily obtained the best Accuracy and F1-macro metrics on both ViMMRC 2.0 and ViRCSoSciD datasets. Moreover, on the ViMMRC 2.0 dataset, our proposed ViMultiChoice has significantly better scores than multi-stage baselines reported from the study Luu et al. (2023). These results strongly highlights the effectiveness of the ViWordFormer module for modeling phrasal lexemes of Vietnamese texts, which enhance the reading and understanding ability of the original OCN for multiple-choice task.
@@ 8 Results Analysis of ViMultiChoice
Model Literature History Geography
Civic
Education
@@
Table 7: F1-macro of ViMultiChoice and baselines for multiple-choice on each subject of the ViRCSoSciD dataset. w/o Expl. stands for ViMultiChoice without Explanation Generator.
On the ViRCSoSciD dataset, ViMultiChoice attains the highest macro F1 score across all Vietnamese high school grade levels. It substantially outperforms ViT5 Encoder and OCN on Grade 10 and Grade 12 questions, indicating the effectiveness of the proposed approach. Subject-level analysis (Table 7) further shows that ViMultiChoice achieves the best macro F1 in most subjects, particularly Civic Education. With the Explanation Generator, the model improves performance across nearly all subjects, except History, where a slight decrease is observed. Overall, ViMultiChoice under both single-task and multi-task training settings demonstrates clear improvements over all baseline models.
@@ 9 Conclusion
Appendix A Appendix: Details of the Option Inference Module
# Retriever P@10 R@10 P@15 R@15 P@20 R@20 P@30 R@30 P@40 R@40 P@50 R@50
1 BM25 28.04 03.75 24.29 04.77 21.89 05.66 18.68 07.11 16.59 08.33 15.06 09.38
2 MiniLML12 29.42 03.92 26.57 05.23 24.61 06.39 21.85 08.39 19.95 10.13 18.50 11.67
3 MultiE5 38.52 05.25 34.07 06.86 31.05 08.24 26.94 10.56 24.17 12.49 22.10 14.16
4 VietBGE 42.77 05.78 38.12 07.60 34.75 09.12 30.16 11.69 26.96 13.79 24.53 15.55
5 VietBGEMultiE5 42.53 05.75 37.92 07.57 34.56 09.10 30.12 11.71 26.92 13.80 24.53 15.58
$f_{O_{k}}^{Q}=[f_{Q};f_{O_{k}}]\in\mathbb{R}^{n_{Qk}\times d},k\in\{1,2,3,4\}$
where $[;]$ denotes the concatenation operator along the last dimension and $n_{Qk}=n_{Q}+n_{k}$. First, we compare any option $O_{k}$ with another option $O_{l}(l\neq k)$ via the trilinear attention mechanism Ran et al. [2019]:
$f_{O_{k,l}}^{Q}=Attn(f_{O_{k}}^{Q},f_{O_{l}}^{Q})f_{O_{l}}^{Q}$ (13)
where $W_{C}\in\mathbb{R}^{4d\times d}$ and $b_{C}\in\mathbb{R}^{d}$.
$f^{O}_{k}=g_{k}\odot f_{O_{k}}^{Q}+(1-g_{k})\odot\bar{f}^{Q}_{O_{k}}$ (18)
$A^{O}_{k}=Attn(f_{k}^{O},f_{P})\in\mathbb{R}^{n_{Qk}\times n_{P}}$ (20)
$A^{P}_{k}=Attn(f_{P},f_{k}^{O})\in\mathbb{R}^{n_{P}\times n_{Qk}}$ (21)
$\hat{f}^{O}_{k}=A^{O}_{k}[f_{P};A^{P}_{k}f^{O}_{k}]\in\mathbb{R}^{n_{Qk}\times 2d}$ (22)
where
and
Appendix B Appendix: Experiments of Retrieval Methods
Appendix C Appendix: Detailed Configurations of LLMs
@@ 8 Mistral-Large mistral-large-2 123B
Model Literature History Geography Civic Education
Accuracy F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy F1-macro
Flan-T5 31.08 25.27 25.42 20.31 24.52 17.33 24.73 19.97
Vicuna 24.32 18.73 36.82 24.50 25.48 15.16 37.36 25.96
Ura-Llama 35.13 26.19 30.17 28.04 28.36 19.78 36.56 28.00
Vina-Llama 40.54 26.29 40.38 39.51 31.97 29.26 44.35 35.38
Qwen2 37.84 18.18 52.26 18.99 51.20 22.81 62.10 31.44
Gemini-Flash 66.22 66.22 71.26 57.07 68.51 54.83 75.81 60.85
Mistral-Large 52.70 52.41 60.09 60.04 63.70 63.40 72.58 72.59
Table 10: Results (%) of LLMs on the multiple-choice task of the ViRCSoSciD dataset across different subjects. Bold numbers indicate the best results.
Model Literature History Geography Civic Education
BLEU-4 ROUGE-L BLEU-4 ROUGE-L BLEU-4 ROUGE-L BLEU-4 ROUGE-L
Flan-T5 0.25 8.64 0.18 6.64 0.40 11.14 0.24 8.17
Vicuna 9.11 32.24 9.26 32.61 9.65 30.89 17.71 41.59
Ura-Llama 11.88 36.34 7.85 32.10 11.09 34.35 19.35 41.96
Vina-Llama 11.25 34.88 12.19 36.70 12.07 34.32 18.19 39.89
Qwen2 9.94 31.77 11.86 35.71 11.40 31.74 13.51 31.61
Gemini-Flash 8.79 42.17 10.57 38.74 14.34 41.34 17.16 44.18
Mistral-Large 13.35 41.97 13.04 39.54 16.04 41.60 21.76 46.03
Table 11: Results (%) of LLMs on the explanation generation task of the ViRCSoSciD dataset across different subjects. Bold numbers indicate the best results.
@@
Hy chn p n ng (A, B, C hoc D),
sau  cho li gii thch ngn gn.
p n:
Prompt for Zero-shot Vincuna
Bn l mt tr l thng minh. Hy tr
li ngi dng mt cch chnh xc bng
cch chn p n ng, ghi ngn gn A,
B, C hay D, sau  cho li gii thch
ngn gn.
@@
    Hy chn p n ng, ghi ngn gn
    A, B, C hay D, sau  cho li gii
    thch ngn gn.
Appendix D Appendix: More Results of LLMs on the ViRCSoSciD Dataset

### v2

@@ 5.3 Configuration
All baseline models and the proposed method ViMultiChoice were trained using the Adam optimizer [10]. For baseline models based on static word embeddings, we employed pretrained word embedding PhoW2V [18] at syllable level, while the remaining baselines utilized the pretrained language model ViT5 [25] to encode input sequences. Due to the maximum input length constraint of mBERT (512 tokens), the lengths of the passage, question, and answer options were capped at 400, 80, and 20 tokens, respectively, to preserve essential information for the MCRC task. All methods were trained with a learning rate of $5\times 10^{-5}$, a batch size of 32 on an NVIDIA A100 GPU. The interruption of the training process is evolve if there is no improvement in F1-macro after 10 consecutive epochs.
Moreover, we deployed an information retriever to obtain the discrete sentences for training baselines and ViMultiChoice method on the ViRCSoSciD dataset. Through our experiments, we found that VietBGE, the pretrained BGE-M3 [4] specially fine-tuned for Vietnamese is the most effective retriever.
@@ 6.1 Results of LLMs on the ViRCSoSciD Dataset
Table 1 denotes the results of LLMs for the multiple choice task while Table 2 indicates the results of LLMs for explanation task on our ViRCSoSciD dataset. Mitral-Large obtained the best performance on F1-macro on multiple-choice task and BLEU-4 as well as ROUGE-L on explanation generator. Gemini-Flash achieved the second position on both tasks, especially it has the best accuracy score on giving options for multiple-choice.
@@
However, the performance of other LLMs is left far away from the two best LLMs although they were pretrained on extremely huge corpora. Moreover, Mistral-Large still gives a slightly low ROUGE-L score for explanation task. These numbers highlights the challenging of our dataset as well as the need of developing more effective LLM particularly for Vietnamese and other low-resource languages.
Model Literature History Geography Civic Education
Flan-T5 25.27 20.31 17.33 19.97
Vicuna 18.73 24.50 15.16 25.96
Ura-Llama 26.19 28.04 19.78 28.00
Vina-Llama 26.29 39.51 29.26 35.38
Qwen2 18.18 18.99 22.81 31.44
Gemini-Flash 66.22 57.07 54.83 60.85
Mistral-Large 52.41 60.04 63.40 72.59
Table 3: F1-macro (%) of LLMs on the multiple-choice task of the ViRCSoSciD dataset across different subjects. Bold numbers indicate the best results.
Model Literature History Geography Civic Education
B-4 R-L B-4 R-L B-4 R-L B-4 R-L
Flan-T5 0.25 8.64 0.18 6.64 0.40 11.14 0.24 8.17
Vicuna 9.11 32.24 9.26 32.61 9.65 30.89 17.71 41.59
Ura-Llama 11.88 36.34 7.85 32.10 11.09 34.35 19.35 41.96
Vina-Llama 11.25 34.88 12.19 36.70 12.07 34.32 18.19 39.89
Qwen2 9.94 31.77 11.86 35.71 11.40 31.74 13.51 31.61
Gemini-Flash 8.79 42.17 10.57 38.74 14.34 41.34 17.16 44.18
Mistral-Large 13.35 41.97 13.04 39.54 16.04 41.60 21.76 46.03
Table 4: BLEU-4 (B-4) (%) and ROUGE-L (R-L) (%) of LLMs on the explanation generation task of the ViRCSoSciD dataset across different subjects. Bold numbers indicate the best results.
6.2 Results of ViMultiChoice
6.2.1 Results of Retrieval Methods
Retriever P@10 R@10 P@15 R@15 P@20 R@20 P@30 R@30
BM25 28.04 03.75 24.29 04.77 21.89 05.66 18.68 07.11
MiniLML12 29.42 03.92 26.57 05.23 24.61 06.39 21.85 08.39
MultiE5 38.52 05.25 34.07 06.86 31.05 08.24 26.94 10.56
[…]
6.2.2 Results on the MCRC tasks
@@
Especially, in comparison to the original OCN, our ViMultiChoice can easily obtained the best Accuracy and F1-macro metrics on both ViMMRC 2.0 and ViRCSoSciD datasets. Moreover, on the ViMMRC 2.0 dataset, our proposed ViMultiChoice has significantly better scores than multi-stage baselines reported from the study [15]. These results strongly highlights the effectiveness of the ViWordFormer module for modeling phrasal lexemes of Vietnamese texts, which enhance the reading and understanding ability of the original OCN for multiple-choice task.
@@ 7 Ablation study of ViMultiChoice
The novelty of our ViMultiChoice is the introduction of the ViWordFormer module for improving Vietnamese text representation and the Explanation Generator for providing explanation of the decision. To highlight the effectiveness of the ViWordFormer module as well as the advantage of having the Explanation Generator, we conducted ablation studies and provided the evaluation in Table 7.
Dataset VWF Type Multiple Choice Explanation
Accuracy F1-macro BLEU-4 ROUGE-L
ViRCSoSciD Single Task 61.81 61.79 - -
62.98 63.03 - -
ViMMRC 2.0 56.27 56.29 - -
60.32 60.23 - -
ViRCSoSciD Multitask 62.51 62.51 20.23 44.71
64.38 64.38 34.85 56.78
Table 7: Ablation study of ViWordFormer (VWF) module for ViMultiChoice on the ViRCSoSciD and ViMMRC 2.0 datasets.
@@ 8 Results Analysis of ViMultiChoice
Model Literature History Geography Civic Education
@@
On the ViRCSoSciD dataset, ViMultiChoice attains the highest macro F1 score across all Vietnamese high school grade levels (Table 9). It substantially outperforms ViT5 Encoder and OCN on Grade 10 and Grade 12 questions, indicating the effectiveness of the proposed approach. Subject-level analysis (Table 10) further shows that ViMultiChoice achieves the best macro F1 in most subjects, particularly Civic Education. With the Explanation Generator, the model improves performance across nearly all subjects, except History, where a slight decrease is observed. Overall, ViMultiChoice under both single-task and multi-task training settings demonstrates clear improvements over all baseline models.
@@ 9 Conclusion
Appendix A Detailed Configurations of LLMs
@@ 8 Mistral-Large mistral-large-2 123B
Hãy chọn đáp án đúng (A, B, C hoặc D), sau đó cho lời giải
thích ngắn gọn.
Đáp án:
Prompt for Zero-shot Vicuna
Bạn là một trợ lý thông minh. Hãy trả lời người dùng một
cách chính xác bằng cách chọn đáp án đúng, ghi ngắn gọn
A, B, C hay D, sau đó cho lời giải thích ngắn gọn.
@@
    Hãy chọn đáp án đúng, ghi ngắn gọn A, B, C hay D, sau
    đó cho lời giải thích ngắn gọn.
Appendix B Visualization of Results
We provide some results of ViMultiChoice for qualitative demonstration. In particular, ViMultiChoice and determine correctly the topic and subject of the given question (Figure 7). This is indicated in the explanation it generated. This advantage help this model give the correct answers and the comprehensive explanation for its decision. However, on questions having confusing or closely similar information (Figure 8), ViMultiChoice tends to have incorrect decision even though it gave reasonable explanations. The further improvement on ViMultiChoice should rely on improving the comparison of options against each other to enhance the ability of determining correctly answer among confusing ones.
Figure 7: Examples for the correct option decision of ViMultiChoice.
Figure 8: Examples for the wrong option decision of the ViMultiChoice.